# In the Picture: Medical Imaging Datasets, Artifacts, and their Living Review

### v1

@@ Abstract.
Disclaimer: this is a working paper, and represents research in progress. We welcome contributions from the community, for comments or questions please email us at amji@itu.dk and vech@itu.dk.
Datasets play a critical role in medical imaging research, yet issues such as label quality, shortcuts, and metadata are often overlooked. This lack of attention may harm the generalizability of algorithms and, consequently, negatively impact patient outcomes. While existing medical imaging literature reviews mostly focus on machine learning (ML) methods, with only a few focusing on datasets for specific applications, these reviews remain static – they are published once and not updated thereafter. This fails to account for emerging evidence, such as biases, shortcuts, and additional annotations that other researchers may contribute after the dataset is published. We refer to these newly discovered findings of datasets as research artifacts. To address this gap, we propose a living review that continuously tracks public datasets and their associated research artifacts across multiple medical imaging applications. Our approach includes a framework for the living review to monitor data documentation artifacts, and an SQL database to visualize the citation relationships between research artifact and dataset. Lastly, we discuss key considerations for creating medical imaging datasets, review best practices for data annotation, discuss the significance of shortcuts and demographic diversity, and emphasize the importance of managing datasets throughout their entire lifecycle. Our demo is publicly available at http://130.226.140.142.
copyright: none
[…]
Manuscript $\xpatchcmd\ps@standardpagestyleManuscript submitted to ACM$ \@ACM@manuscriptfalse
@@ 1. Introduction
  • We present a living review (publications and database, see Fig. 2) - a framework for enhancing dataset metadata through research artifacts, facilitating the discovery of emerging information about medical imaging datasets.
@@ 2. Proposal: a living review of medical imaging datasets
Traditional systematic reviews of medical imaging datasets, such as (Daneshjou et al., 2021; Wen et al., 2022; Khan et al., 2021; Li et al., 2023), do not capture this evolution. We argue that we need a living review to keep track of novel open datasets, as well as emerging connections and issues in existing datasets. Living systematic reviews are a more recent development in meta-research, but they are crucial for rapidly evolving topics, such as for example evidence around COVID in recent years (Wynants et al., 2020). Given the rapid developments in ML, it would be advantageous to adopt a similar framework for datasets. Various best practices for living reviews already exist in other fields, for example (Schmidt et al., 2021, 2023).
@@
Our living review consists of three parts, as illustrated in Figure 2:
@@
The SQL database consists of three tables (see Fig. C2): datasets, papers documenting their research artifacts, and the dataset-artifact relations. Our interactive demo http://130.226.140.142 (built with PostgreSQL, Python, Streamlit and the st_link_analysis package), supports dynamic dataset-artifact exploration. In the demo, we present 24 research artifacts related to 16 datasets, which include additional annotations, evidence of shortcuts, and derived datasets, focused on two applications: skin lesions and chest X-rays.
@@ 3. Considerations for creating a dataset
With (smaller) public datasets, a common solution is to inject additional knowledge from another source into the dataset, such as domain knowledge provided by experts, or using data or representations from a different source. Transfer learning (Pan and Yang, 2010; Cheplygina et al., 2019) is therefore often used in medical imaging. A common approach is to fine-tune models pretrained on ImageNet (Russakovsky et al., 2015), although recent results show this strategy is more sensitive to shortcuts (Juodelyte et al., 2024) than if training on RadImageNet (Mei et al., 2022), a recent dataset with a million images from different radiological modalities.
@@ 5. Inside a dataset: demographics
The potential of overfitting to established benchmark datasets is a long-standing debate in the ML community (Liao et al., 2021; Gossmann et al., 2018; Hosseini et al., 2020). While the past few years have seen a surge in public datasets, they typically lack demographic metadata about the subjects. This leads to a host of problems. If the data does not come with demographic information, researchers cannot assess whether the datasets and models trained on them have demographic biases. As a result, not only the resulting algorithms, but also what we learn about ML development, can be tainted by demographic biases without our knowledge. Equally important, research on algorithmic bias and fairness, which needs demographic metadata, has access to very few datasets. As a result, large amounts of research resources have been dedicated to a very small set of case studies, whose particularities – for better and for worse – drive the progress of the research field. Here we present four case studies, illustrated in Figure 3, further clinically relevant details in Section B.
@@
Chest X-rays are the most commonly performed radiologic examinations worldwide (UNSCEAR et al., 2000), requiring significant expertise for accurate and meaningful interpretation (Van Ginneken et al., 2001). ML revolutionized chest X-ray diagnosis, highlighted by the release of NIH-CXR14 dataset (Wang et al., 2017) and CheXNet’s model claiming radiologist-level pneumonia detection (Rajpurkar et al., 2017). However, these claims have been criticized for relying on shortcuts (Oakden-Rayner et al., 2020; Jiménez-Sánchez et al., 2023), and for low inter-rater agreement (Damgaard et al., 2023). Subsequently, additional datasets such as CheXpert (Irvin et al., 2019), MIMIC-CXR (Johnson et al., 2019), and PadChest (Bustos et al., 2020) have become widely used in the research community.
@@
The few available datasets are the Fetal Planes DB (Burgos-Artizzu et al., 2020) and data from the HC18 (van den Heuvel et al., 2018), FH-PS-AOP (Jieyun and ZhanHong, 2023), and ACOUSLIC-AI (Sappia, 2024) challenges, with none, to date, providing demographic information. More diverse datasets with detailed demographics, including BMI and ethnicity, image quality information, and expertise of the clinicians, are needed to develop fairer models and improve prenatal screening. This is true for low-, middle- and high-resource countries. In low and middle-income resource settings with poorer image quality due to portable devices this is crucial to reduce maternal complications and fetal mortality ((2024), WHO). An upcoming study in Europe shows that ML using imaging is more robust at predicting birth weight compared to measurements extracted from traditional ultrasound procedures, standards and prediction charts. Future studies should therefore look with care at how both image quality and study design affect performance across groups – for both ML and more traditional predictive approaches.
@@
MRI is the third most commonly performed imaging modality after CT and X-rays. Its superior soft tissue contrast enables detailed visualization of brain anatomy, making it ideal for detecting abnormalities, and most public MRI datasets for ML research focus on the brain (Dishner et al., 2024). Key application areas for brain MRI and notable datasets, include neurodegenerative diseases (OASIS (Marcus et al., 2007, 2010; LaMontagne et al., 2019; Koenig et al., 2020), brain cancer (BraTS (Baid et al., 2021), LUMIERE (Suter et al., 2022), Ocana (Ocaña-Tienda et al., 2023), and TCGA-GBM (Scarpace et al., 2016)), and stroke, (ISLES 2022 (Hernandez Petzsche et al., 2022), ATLAS v2.0 (Liew et al., 2022), among others).
@@ 6. Inside a dataset: shortcuts
Shortcuts can arise from spurious correlations (e.g. noise patterns or differences in intensity distributions across various scanners), demographic attributes, and their interactions. The Medical Imaging Contextualized Confounder Taxonomy (MICCAT) (Juodelyte et al., 2024) (see Fig. A1) helps identify the shortcuts’ origin and mitigation strategies. MICCAT extends beyond traditional demographic attributes to include a broader set of confounders that are domain- and context-specific confounders at both patient and environment levels.
@@ 8. Discussion and conclusions
Although more data-centric initiatives, such as the Datasets and Benchmarks Track at NeurIPS, are emerging, it is clear that more attention to data work is needed. Perhaps in this context it is worth mentioning that despite the importance of the topic and a diverse group of researchers, our own workshop proposal was initially rejected by two different venues. However, given recent progress, we remain hopeful data work will be more addressed and better acknowledged in the future, and we are in the process of applying for funding for related research and networking events, to ensure the continuity of this work. We want to promote a cultural shift toward the responsible use of high quality datasets, created through collaboration with relevant stakeholders or clinical experts. We therefore invite everyone involved in and affected by healthcare ML to contribute to our living review and other dataset efforts.
9. Statements
9.1. Ethical considerations statement
@@
9.2. Adverse impact statement
@@
Acknowledgments
This project has received funding from the Independent Research Council Denmark (DFF) Inge Lehmann 1134-00017B. We extend our gratitude to the speakers and participants of the “Datasets through the Looking-Glass” webinar and the “In the Picture: Medical Imaging Datasets” workshop, which have helped to shape this research. The workshop received funding from the Danish Data Science Academy (DDSA). EF gratefully acknowledges the support of the Google Award for Inclusion Research (AIR) Program. AH is employed at Steno Diabetes Center Aarhus that is partly funded by a donation from the Novo Nordisk Foundation. AH is supported by a Data Science Emerging Investigator grant by the Novo Nordisk Foundation (NNF22OC0076725). TR was supported by a scholarship from the Hanns Seidel Foundation with funds from the Federal Ministry of Education and Research Germany (BMBF). DW received funding from the Jill and Herbert Hunt Scholarship, University of Oxford. MAZ is funded by TRAIN (ANR-22-FAI1-0003-02). We thank Freepick for the icons in Figure 2.
@@ Appendix A Shortcuts
Machine learning models are prone to rely on spurious correlations to make predictions, which are usually easier to detect than the genuine disease patterns, as explained in Section 6. Such shortcuts can be well-localized objects in the image – e.g., mechanical ventilation tubes or pacemakers, which are often present in patients with certain diseases– or more global ones, like specific noise patterns or intensity distributions associated to a certain acquisition setting or device brand. Demographic attributes like sex/gender, age or ethnicity can also lead to shortcut learning, disproportionately impacting historically underserved subgroups (Banerjee et al., 2023), especially when datasets are highly imbalanced. To categorize the different types of shortcuts, we adopt the Medical Imaging Contextualized Confounder Taxonomy (MICCAT) (Juodelyte et al., 2024), see Fig. A1, which we believe can be useful to understand how spurious correlations arise and how they can be mitigated. MICCAT extends beyond traditional demographic attributes, such as sex/gender, age, and ethnicity, to include a broader set of confounders that are domain- and context-specific. These confounders encompass patient-level and environment-level factors.
@@
Figure A1. MICCAT: Medical Imaging Contextualized Confounder Taxonomy. Reproduced from (Juodelyte et al., 2024) with permission of the authors.

### v2

@@ Abstract.
Datasets play a critical role in medical imaging research, yet issues such as label quality, shortcuts, and metadata are often overlooked. This lack of attention may harm the generalizability of algorithms and, consequently, negatively impact patient outcomes. While existing medical imaging literature reviews mostly focus on machine learning (ML) methods, with only a few focusing on datasets for specific applications, these reviews remain static – they are published once and not updated thereafter. This fails to account for emerging evidence, such as biases, shortcuts, and additional annotations that other researchers may contribute after the dataset is published. We refer to these newly discovered findings of datasets as research artifacts. To address this gap, we propose a living review that continuously tracks public datasets and their associated research artifacts across multiple medical imaging applications. Our approach includes a framework for the living review to monitor data documentation artifacts, and an SQL database to visualize the citation relationships between research artifact and dataset. Lastly, we discuss key considerations for creating medical imaging datasets, review best practices for data annotation, discuss the significance of shortcuts and demographic diversity, and emphasize the importance of managing datasets throughout their entire lifecycle. Our demo is publicly available at http://inthepicture.itu.dk/.
open data, data governance, healthcare, medical imaging, shortcuts, bias, research artifacts, living review
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732035isbn: 979-8-4007-1482-5/2025/06ccs: General and reference Surveys and overviewsccs: General and reference Evaluationccs: Applied computing Life and medical sciencesccs: Computing methodologies Machine learning
@@ 1. Introduction
  • We present a proof of concept for a living review (publications and database, see Fig. 2) - a framework for enhancing dataset metadata through research artifacts, facilitating the discovery of emerging information about medical imaging datasets.
@@
[…]
This figures summarizes the structure of this manuscript highlighting the benefits of our proposed living review for medical imaging datasets. It provides an overview of considerations for creating a dataset (Sec. 3), data annotation practices and quality (Sec. 4), patient demographics (Sec. 5) and shortcuts (Sec. 6), and dataset lifecycle (Sec. 7).
@@ 2. Proposal: a living review of medical imaging datasets
Traditional systematic reviews of medical imaging datasets, such as (Daneshjou et al., 2021; Wen et al., 2022; Khan et al., 2021; Li et al., 2023), do not capture this evolution. We argue that we need a living review to keep track of novel open datasets, as well as emerging connections and issues in existing datasets. Living systematic reviews are a more recent development in meta-research, but they are crucial for rapidly evolving topics. Given the rapid developments in ML, it would be advantageous to adopt a similar framework for datasets. Our proposed work is different from these efforts, since (Wynants et al., 2020) focuses on COVID-19 findings, not specifically on medical image datasets. (Schmidt et al., 2021, 2023) focus on data extraction methods to aid doing systematic reviews, of these only (Schmidt et al., 2021) is a living review itself, which could provide inspiration for how to carry out our plans.
@@
Our proposal for the living review consists of three parts, as illustrated in Fig. 2:
@@
[…]
Our living review consists of three components: (1) an overarching living review paper, (2) documentation of research artifacts via publications on Zenodo linked by living review paper, and (3) a SQL database for exploring the links between the datasets and the research artifacts.
@@
The SQL database consists of three tables (see Fig. C2): datasets, papers documenting their research artifacts, and the dataset-artifact relations. Our interactive demo http://130.226.140.142 (built with PostgreSQL, Python, Streamlit and the st_link_analysis package), supports dynamic dataset-artifact exploration. For additional technical details, see Section C. In the demo, we present 24 research artifacts related to 16 datasets, which include additional annotations, evidence of shortcuts, and derived datasets, focused on two applications: skin lesions and chest X-rays.
@@ 3. Considerations for creating a dataset
With (smaller) public datasets, a common solution is to inject additional knowledge from another source into the dataset, such as domain knowledge provided by experts, or using data or representations from a different source. Transfer learning (Pan and Yang, 2010; Cheplygina et al., 2019) is therefore often used in medical imaging. A common approach is to fine-tune models pretrained on ImageNet (Russakovsky et al., 2015), although recent results show this strategy is more sensitive to shortcuts (Juodelyte et al., 2025) than if training on RadImageNet (Mei et al., 2022), a recent dataset with a million images from different radiological modalities.
@@ 5. Inside a dataset: demographics
[…]
Case studies: (a) A normal posteroanterior chest radiograph of a patient with no visible signs of injury, showing letters that indicate left or right. (b) A malignant melanoma dermoscopic image from ISIC dataset, featuring dark corners and areas of intense brightness. (c) A cross-sectional T1-weighted MRI of a healthy human brain. (d) A fetal ultrasound image displaying the head of the fetus showing information about the head diameter and gestational age.
@@
The potential of overfitting to established benchmark datasets is a long-standing debate in the ML community (Liao et al., 2021; Gossmann et al., 2018; Hosseini et al., 2020). While the past few years have seen a surge in public datasets, they typically lack demographic metadata about the subjects. This leads to a host of problems. If the data does not come with demographic information, researchers cannot assess whether the datasets and models trained on them have demographic biases. As a result, not only the resulting algorithms, but also what we learn about ML development, can be tainted by demographic biases without our knowledge. Equally important, research on algorithmic bias and fairness, which needs demographic metadata, has access to very few datasets. As a result, large amounts of research resources have been dedicated to a very small set of case studies, whose particularities – for better and for worse – drive the progress of the research field. Here we present four case studies, illustrated in Fig. 3, further clinically relevant details in Section B.
@@
Chest X-rays are the most commonly performed radiologic examinations worldwide (UNSCEAR et al., 2000), requiring significant expertise for accurate and meaningful interpretation (Van Ginneken et al., 2001). ML revolutionized chest X-ray diagnosis, highlighted by the release of NIH-CXR14 dataset (Wang et al., 2017) and CheXNet’s model claiming radiologist-level pneumonia detection (Rajpurkar, 2017). However, these claims have been criticized for relying on shortcuts (Oakden-Rayner et al., 2020; Jiménez-Sánchez et al., 2023), and for low inter-rater agreement (Damgaard et al., 2023). Subsequently, additional datasets such as CheXpert (Irvin et al., 2019), MIMIC-CXR (Johnson et al., 2019), and PadChest (Bustos et al., 2020) have become widely used in the research community.
@@
The few available datasets are the Fetal Planes DB (Burgos-Artizzu et al., 2020) and data from the HC18 (van den Heuvel et al., 2018), FH-PS-AOP (Jieyun and ZhanHong, 2023), and ACOUSLIC-AI (Sappia, 2024) challenges, with none, to date, providing demographic information. More diverse datasets with detailed demographics, including BMI and ethnicity, image quality information, and expertise of the clinicians, are needed to develop fairer models and improve prenatal screening. This is true for low-, middle- and high-resource countries. In low and middle-income resource settings with poorer image quality due to portable devices this is crucial to reduce maternal complications and fetal mortality ((2025), WHO). A recent Danish study (Mikolaj et al., 2025) shows that across demographic subgroups, deep learning algorithms improve birth weight estimates from fetal ultrasound compared to clinical standard measurements extracted from those same ultrasound images – potentially because the clinical standard measurements are unable to use additional image content to make up for suboptimal ultrasound planes. Future studies should therefore look with care at how both image quality and study design affect performance across groups – for both ML and more traditional predictive approaches.
@@
MRI is the third most commonly performed imaging modality after CT and X-rays. Its superior soft tissue contrast enables detailed visualization of brain anatomy, making it ideal for detecting abnormalities, and most public MRI datasets for ML research focus on the brain (Dishner et al., 2024). Key application areas for brain MRI and notable datasets, include neurodegenerative diseases (OASIS (Marcus et al., 2007, 2010; LaMontagne et al., 2019; Koenig et al., 2020)), brain cancer (BraTS (Baid et al., 2021), LUMIERE (Suter et al., 2022), Ocana (Ocaña-Tienda et al., 2023), and TCGA-GBM (Scarpace et al., 2016)), and stroke, (ISLES 2022 (Hernandez Petzsche et al., 2022), ATLAS v2.0 (Liew et al., 2022), among others).
@@ 6. Inside a dataset: shortcuts
Shortcuts can arise from spurious correlations (e.g. noise patterns or differences in intensity distributions across various scanners), demographic attributes, and their interactions. The Medical Imaging Contextualized Confounder Taxonomy (MICCAT) (Juodelyte et al., 2025) (see Fig. A1) helps identify the shortcuts’ origin and mitigation strategies. MICCAT extends beyond traditional demographic attributes to include a broader set of confounders that are domain- and context-specific confounders at both patient and environment levels.
@@ 8. Discussion and conclusions
Although more data-centric initiatives, such as the Datasets and Benchmarks Track at NeurIPS, are emerging, it is clear that more attention to data work is needed. Perhaps in this context it is worth mentioning that despite the importance of the topic and a diverse group of researchers, our own workshop proposal was initially rejected by two different venues. However, given recent progress, we remain hopeful data work will be more addressed and better acknowledged in the future, and we are in the process of applying for funding for related research and networking events, to ensure the continuity of this work.
Looking ahead, established peer-review processes in conferences could be adapted to support the publication and recognition of data artifacts, assigning them DOIs similar to standard papers. Conferences like Medical Imaging with Deep Learning (MIDL) [^2] https://2025.midl.io/ could invite data artifacts as papers for a special track, similar to the Datasets and Benchmarks track at NeurIPS. However, a more scalable approach would involve multiple conferences adopting a “rolling” review style, as has been adopted in the natural language processing community (Review, 2025). Another interesting option would be for conferences to invite collaborative contributions to selected datasets, similar to medical imaging competitions like those hosted at the International Conference on Medical Image Computing and Computer Assisted Interventions (MICCAI) [^3] https://conferences.miccai.org/2025/en/challenges.asp. In the existing scenario, the organizers select a dataset, competition participants develop algorithms, and after benchmarking, both organizers and many participants typically co-author a paper about the results. Contributions to a datasheet would, of course, not be trivial to “benchmark”; however, there would not need to be a “winner-takes-all” mentality. If different teams highlight different aspects of the datasets, their contributions could simply be combined.
With our work, we want to promote a cultural shift toward the responsible use of high quality datasets, created through collaboration with relevant stakeholders or clinical experts. We therefore invite everyone involved in and affected by healthcare ML to contribute to our living review and other dataset efforts.
Ethical considerations statement
@@
Adverse impact statement
@@
Author Contributions
Authors, except for the first and last, are listed in alphabetical order. Contributions follow the CRediT author statement.
Conceptualization, Methodology, Writing - Original Draft & Final Draft: Veronika Cheplygina, Amelia Jiménez-Sánchez.
Data Curation, Visualization, Software: Amelia Jiménez-Sánchez.
Supervision, Funding Acquisition: Veronika Cheplygina.
Writing - Review & Editing: all authors contributed, with individual section contributions listed below in alphabetical order:
  • Introduction: Veronika Cheplygina, Amelia Jiménez-Sánchez, Maria A. Zuluaga
  • Proposal: a living review of medical imaging datasets: Veronika Cheplygina, Amelia Jiménez-Sánchez
  • Considerations for creating a dataset: Veronika Cheplygina, Camila González, Steff Groefsema, Amelia Jiménez-Sánchez, Leo Joskowicz, Melih Kandemir, Hubert Dariusz Zając.
  • Data annotation practices and quality: Veronika Cheplygina, Alessa Hering, Leo Joskowicz, Thijs Kooi, Tim Rädsch.
  • Inside a dataset: demographics: Víctor M. Campello, Aasa Feragen, Jorge del Pozo Lérida, Andre Pacheco, Mauricio Reyes, Nina Weng, Jack Junchi Xu.
  • Inside a dataset: shortcuts: Enzo Ferrante, Judy Wawira Gichoya, Dovile Juodelyte.
  • Data lifecycle: Natalia-Rozalia Avlona, Sarah de Boer, Melanie Ganz-Benjaminsen, Bram van Ginneken, Amelia Jiménez-Sánchez, Théo Sourget.
  • Discussion: Veronika Cheplygina, Adam Hulman, Amelia Jiménez-Sánchez, Livie Yumeng Li, David Wen.
Acknowledgements.
This project has received funding from the Independent Research Council Denmark (DFF) Inge Lehmann 1134-00017B. The “In the Picture: Medical Imaging Datasets” workshop was funded by DFF and the Danish Data Science Academy (DDSA) with Grant ID 2024-2342. We extend our gratitude to the speakers and participants of the “Datasets through the Looking-Glass” webinar, who have helped to shape this research. EF gratefully acknowledges the support of the Google Award for Inclusion Research (AIR) Program. AJS was financed by the DFF grant MMC. AH is employed at Steno Diabetes Center Aarhus that is partly funded by a donation from the Novo Nordisk Foundation. AH is supported by a Data Science Emerging Investigator grant by the Novo Nordisk Foundation (NNF22OC0076725). TR was supported by a scholarship from the Hanns Seidel Foundation with funds from the Federal Ministry of Education and Research Germany (BMBF). DW received funding from the Jill and Herbert Hunt Scholarship, University of Oxford. MAZ is funded by TRAIN (ANR-22-FAI1-0003-02). We thank Freepick for the icons in Fig. 2.
@@ Appendix A Shortcuts
Machine learning models are prone to rely on spurious correlations to make predictions, which are usually easier to detect than the genuine disease patterns, as explained in Section 6. Such shortcuts can be well-localized objects in the image – e.g., mechanical ventilation tubes or pacemakers, which are often present in patients with certain diseases– or more global ones, like specific noise patterns or intensity distributions associated to a certain acquisition setting or device brand. Demographic attributes like sex/gender, age or ethnicity can also lead to shortcut learning, disproportionately impacting historically underserved subgroups (Banerjee et al., 2023), especially when datasets are highly imbalanced. To categorize the different types of shortcuts, we adopt the Medical Imaging Contextualized Confounder Taxonomy (MICCAT) (Juodelyte et al., 2025), see Fig. A1, which we believe can be useful to understand how spurious correlations arise and how they can be mitigated. MICCAT extends beyond traditional demographic attributes, such as sex/gender, age, and ethnicity, to include a broader set of confounders that are domain- and context-specific. These confounders encompass patient-level and environment-level factors.
@@
Figure A1. MICCAT: Medical Imaging Contextualized Confounder Taxonomy. Reproduced from (Juodelyte et al., 2025) with permission of the authors.
[…]
MICCAT: Medical Imaging Contextualized Confounder Taxonomy. The patient level can be further classified into anatomical confounders and demographic attributes. The environment level can be further classified into imaging confounders and external confounders (physical or virtual devices).